A rope of 18 metre is used to form a sector of a circle of radius 3.5 meter on a school Playfield. What is the size of the angle of the sector

Found 2 solutions by Alan3354, Cromlix:
Answer by Alan3354(69443)

A rope of 18 metre is used to form a sector of a circle of radius 3.5 meter on a school Playfield. What is the size of the angle of the sector
---------
Angle = 18/3.5 radians =~ 5.14 radians
---------
If the rope is all 3 sides:
2 sides are 3.5m each = 7 meters
The angle = 11/3.5 radians =~ 3.14 radians

Answer by Cromlix(4381)

Hi there,
With radius of sector equaling 3.5 metres
then length of arc must be 11 metres.
18 - 2(3.5) = 11
Angle/360 = Length of Arc/π*2*radius [* = times]
Angle/360 = 11 m/π * 7
(Angle*π* 7)/(11 * 360)
Angle = (11 * 360)/π * 7
Angle = 180.1 degrees. (1 decimal place)
Hope this helps :-)